OpenStudy (anonymous):

If r(x) = 3x − 1 and t(x) = −4x − 1, find (r + t)(7). 9 −9 −49 49

OpenStudy (anonymous):

plug and chug

OpenStudy (anonymous):

r(x) is the same thing as r.... only diff is it is in terms of x [(3x - 1) + (-4x -1) ] * 7

OpenStudy (anonymous):

-49?

OpenStudy (anonymous):

no... it just means add the two functions together...: (r+t)(7) = r(7) + t(7) = (3*7 - 1) + (-4*7 - 1) = (20) + (-29) = -9 consequently, you can also simplify (r+t)(x) an calculate from there: (r+t)(x) = (3x-1) + (-4x-1) = -x - 2 so (r+t)(x) = -(7) - 2 = -9